OpenStudy (anonymous):

Enter the quadratic equation in standard form. 2x^2-12+2x=(x-2)^2+2

OpenStudy (anonymous):

2x^2-12+2x=x^2-4x+4+2 x^2+6x-18=0

OpenStudy (anonymous):

Thank you!